the sum of two numbers is 20. The difference of the two numbers is 8. What are the two numbers??? 16 and 4
Found 2 solutions by Earlsdon, melanie125:
Answer by Earlsdon(6103)

Let x and y be the two numbers.
x+y = 20
x-y = 8 Add these two equations.
-------
2x = 28 Divide both sides by 2.
x = 14
y = 20-x
y = 6
The two numbers are: 14 and 6

Answer by melanie125(2)